The product of two consecutive odd integers is 1 less than 5 times their sum. Find the two integers. Answer in the form of paired points with the lowest of the two integers first

Answer by CubeyThePenguin(3113)

consecutive odd integers: x, (x+2)

x(x+2) = 5(x + x+2) - 1
x^2 + 2x = 5(2x + 2) - 1
x^2 + 2x = 10x + 9
x^2 - 8x - 9 = 0
(x - 9)(x + 1) = 0

The integers could be (-1, 1) or (9, 11).